A. Write an algorithm in pseudocode, which prints all strings of length 6 where the first three characters are any symbol from X={!,@,#,$,%,^}, the 4th and 5th characters are any digit 0,1,…9, and the last character is any digit 1,2,…9.

How many times will the print command be called (e.g. how many 6-character strings will be printed)?

B. Let g(n)=n3 + 2n2 - 5. Determine the big-theta estimate of f. A complete response will include analysis of: k3HhXiTAAAAAElFTkSuQmCC, and qAAAAAElFTkSuQmCC.

C. How many compares of the form fExeT3y45evk4AAAAAElFTkSuQmCC   will the following algorithm perform on a list of length n? Express your answer as a rational expression in terms of n.

for 4Pi0+1u0xuwAAAAASUVORK5CYII= do

         for SEargks2AAAAAElFTkSuQmCC do

                 if fExeT3y45evk4AAAAAElFTkSuQmCC then swap 8GYbT8mpCzFa4AAAAASUVORK5CYII= and 0AZF9BIt7GBqkAAAAASUVORK5CYII=
Answer #1

Sol. Given! prints all strings of length 6. first three characters are any synd for x=!,@,#,#2:1.,1}.. y = {0, 1, 2} Z ={1,2,@ o 2 r for it to length (2) for jki length lyn lr for kal length (3) 6 [print (x(i), y(), alk), x(e)] Given P sendood the prüsn-2 → j =net til i=n-1 ;=$} in total = 0+1+---+(n-3)+(n-2) In-2)(n-2+1) 2 Total = ln-1 (n-3)